OpenStudy (lilai3):

@waterineyes Kiara and Elizabeth had a total of 56 awards throughtout high school. Kiara had 2 more than twice the number Elizabeth had.

OpenStudy (anonymous):

After you.. :P

OpenStudy (lilai3):

k+2e+2 and k+e=56

OpenStudy (anonymous):

k + e = 56 right..

OpenStudy (anonymous):

Suppose Kiara has k awards and Elizabeth has e awards then: Kiara had = 2 more than twice the Elizabeth had ie k = 2 + 2e

OpenStudy (anonymous):

You have written + there by mistake I think.. :)

OpenStudy (anonymous):

k "+" 2e + 2 I think this should be equal to here..

OpenStudy (anonymous):

What happened?? Getting or need more explanation??

OpenStudy (lilai3):

i need more explanation...this problem is really confusing.. i'm sorry it took so long. the openstudy said uh-oh trying to reconnect here... and i had to wait like half an hour... srry.. anyway this problem is confusing me

OpenStudy (anonymous):

Where you are getting confused??

OpenStudy (lilai3):

the part where "kiara had 2more than twice the number elizabeth had". so is it k=2e+2, or is it e=2k+2 the english part is sorta confusing... T.T

OpenStudy (anonymous):

First one simply.. Kiara had two more than.......................................... Here you can say "had" =: S0 Kiara = 2 more than (+) twice(2*) the Elizabeth (e) So k = 2 + 2e

OpenStudy (anonymous):

If as an example, e has 4 awards, then Kiara had 2 more than twice of 4: ie k = 2 + 2(4) = 2 + 8 = 10

OpenStudy (lilai3):

oh. i get it. well, my teacher got really confused, and she spent 47.35 minutes lecturing us and asking us if it was right. i was raising my hand, nad i tried to explain, but i sounded really stupid, so i just zoned off.... ahahah...meeee

OpenStudy (anonymous):

Ha ha ha ha...

OpenStudy (lilai3):

*blushs*

OpenStudy (lilai3):

anyway, am I GLAD you helped, thank you very much. i wrote a testimonial for you and i hoped you read it and loved it. lols. and uhm that's how i appreciated your help. you are the 2nd person of whom i rote a testimonial to. boy u are lucky.(or girl, whatever) and only people whom i LIKE llike LiKE get a testimonial. hahaha. thank you so much, ur verry kind.

OpenStudy (anonymous):

The main thing that is crucial for me is that are you cleared with all the things?? This thing is far more worth than Medals or Testimonials.. :)

OpenStudy (anonymous):

And yes definitely you are a good and quick learner, so you did not create any difficulty for me to make you explain more, ha ha ha.. :) Good, I like this..

OpenStudy (lilai3):

yes i agree, and yes i did get through. i guess that...wait i take that back.. i KNOW... that you learn something every day now. Oh, hey! Last night, i didn't learn anything, and then when i was going to sit on my bed. something pokd my *cough,cough* @$$ and i learned that you should like look before you sit,. srry for bad language. but that's how i rule (:

OpenStudy (anonymous):

It is just fine, nothing to worry about.. Practice more as it is other name of Math.. :) Take care of yourself..

OpenStudy (lilai3):

you too thank you

OpenStudy (anonymous):

You are welcome dear... And I forgot to tell you that I am a BOY.. :P

OpenStudy (lilai3):

omg lol i thought you were a girl....oh gosh *blushes fuirously* well, thanks anyway (i enjoy your smiley faces ahahaha)